# Net Charge Accretion in Magnetized Kerr Black Holes

### v2

@@ IV.3 Critical points and their classification
We conclude that all globally connecting and disconnecting points must correspond to solutions of this set of equations [^5] This system of equations can easily be converted to a set of polynomial equations in all its variables, by replacing the variable $\theta$ with $v\equiv\sin^{2}\theta$ – which is more convenient for analysis.. This fact provides a convenient method for finding the globally (dis)connecting points by numerically solving this set of equations. Mathematically, this set of equations typically has numerous solutions but we restrict ourselves to “physical” solutions, which are real solutions with $r>r_{+},0<\sin^{2}{\theta}\leq 1$ and $b>0$. We refer to these “physical” solutions as critical points [^6] Not to be confused with the notion of “critical trajectory” which corresponds to $b=b_{1}$ mentioned in Subsec. III.2 ., [^7] Interestingly, critical points can be shown to fully correspond to circular trajectories at a (not necessarily equatorial) constant $\theta$. We hope to elaborate on this elsewhere.. We empirically find that there always are at least one and at most four such critical points (for given $0<\alpha<1,\>E>1$ and $\varepsilon\neq 0$) [^8] Practically, we find the critical points by solving the polynomial variant of this set of equations (see footnote 5). To this end, we use Wolfram Mathematica’s command NSolve, which solves this algebraic system easily and with any desired precision. .
@@
Among the relevant points, some satisfy $n_{\mathrm{u},b}>0$, which implies that two allowed regions locally disconnect as $b$ is increased at this point in space; see Fig. 3. We refer to such relevant points as locally disconnecting points. Likewise, relevant points at which $n_{\mathrm{u},b}<0$, will be referred as locally connecting points. Any globally (dis)connecting point must in particular be a locally (dis)connecting point. In particular, note that if there is only one locally disconnecting point, it must correspond to $b_{0}$.
As mentioned above, in principle there might be “energetic rings”. In this case, there must be at least one globally connecting points – and also at least two globally disconnecting ones. As a matter of fact, based on the local analysis of critical points developed above, we find that there are no locally connecting points. This of course implies that there are no globally connecting points, and thus, there are no “energetic rings”. This means that the entire domain $b<b_{0}$ is allowed, and the entire domain $b>b_{0}$ is forbidden.
@@
Using the latter method we find that in the specific case presented in Subsec. V.2 ($\alpha=0.75,E=\sqrt{2}$), in the first four domains there is only a single locally disconnecting point – which must be $b_{0}$. In particular, this is the situation for the specific value $\varepsilon=2$ discussed in further details in Subsec. V.1 (this particular $\varepsilon$ value belongs to the first domain). In the fifth domain, corresponding to $\varepsilon$ values larger than the last transition point $\varepsilon_{\mathrm{last}}\approx 6.8702$, there actually are two locally disconnecting points. The identification of $b_{0}$ in this case is done using a complementary analytical method as we describe in the next subsection. The implementation of the classification process for this last $\varepsilon$ domain is demonstrated in App. D.
@@ IV.4 An analytical upper bound
Considering first the situation at $b=b_{h}$, in App. B we show that at $r=r_{+}$, $n_{\mathrm{u},r}>0$ for all $\theta$ (including $\theta=0$), which implies that $n_{\mathrm{u}}>0$ in a $r>r_{+}$ neighborhood of the event horizon. In other words, particles coming towards the BH with $b=b_{h}$ cannot be absorbed as their presence near the event horizon is forbidden. Because $n_{\mathrm{u}}$ is continuous (particularly in $b$), there must exist a neighborhood of $b=b_{h}$ throughout which the horizon is still surrounded by a forbidden region.
@@ V Charge Accretion (Im)balance
As explained in sections III and IV, the lower bound on the absorption cross section can be found by numerically evaluating $b_{1}$, and the upper bound by finding $b_{0}$ – or alternatively using $b_{h}$. We implemented these methods using Wolfram Mathematica to compute these two bounds. Throughout this section we focus of the specific case of $\alpha=0.75$ and $E=\sqrt{2}$. In the first subsection we specifically consider the value $\mathopen{}\mathclose{{\left|\varepsilon}}\right|=2$ and demonstrate that charge accretion is not balanced at $Q=Q_{\mathrm{w}}$. In the second subsection we extend the analysis to the entire range of $\varepsilon$ and find that charge accretion imbalance persists for all $\mathopen{}\mathclose{{\left|\varepsilon}}\right|$ values greater than a certain, relatively small, parameter $\varepsilon_{\mathrm{cross}}$ to be specified below.
@@ V.1 A specific example of accretion imbalance
In this subsection we focus on the specific case $|\varepsilon|=2$ (with $E=\sqrt{2}$ and $\alpha=0.75$). In this case, following the method described in Subsec. III.2, we first find that for the attracted charge $b_{1}^{-}\approx 4.216$. Then, implementing the methods described in Sec. IV for the repelled charge, we find that at $\varepsilon=2$ there is only a single critical point $(r_{\mathrm{crit}}\approx 1.706,\theta_{\mathrm{crit}}=\frac{\pi}{2},b_{\mathrm{crit}}\approx 2.427)$. A map of the allowed and forbidden regions at this point is presented in Fig. 5. As explained in Subsec. IV.2 (combined with Subsec. IV.3), in such a case, $b_{\mathrm{crit}}$ must correspond to $b_{0}$ – namely, $b_{0}^{+}\approx 2.427$. This is a relevant critical point (namely, it satisfies $\det H<0$ and $n_{\mathrm{u},rr}<0$) which is also locally disconnecting ($n_{\mathrm{u},b}>0$) – as it should.
@@
Figure 5: Map of the allowed and forbidden regions at $b=b_{0}$ for $\alpha=0.75,E=\sqrt{2}$ and $\varepsilon=2$. For this choice of $\varepsilon$ value there is only one critical point (circled in red), which must be $b_{0}$. In this case, $b_{0}\approx 2.4273$.
@@
In this subsection we generalize the analysis and results presented in the previous subsection to the entire $\varepsilon$ axis. As already mentioned above, to find $b_{0}$, we carry this exploration of the $\varepsilon$ axis using two independent methods (both yielding precisely the same results). The first one involves a dense sampling of $\varepsilon$ values along the $\varepsilon$ axis: we scan the range $0<\mathopen{}\mathclose{{\left|\varepsilon}}\right|\leq 10$ with an increment $\Delta\varepsilon=0.01$, and in addition, we scan the larger range $0<\mathopen{}\mathclose{{\left|\varepsilon}}\right|\leq 100$ with $\Delta\varepsilon=0.1$. The second method is more analytical in nature and requires the investigation of only a single representative $\varepsilon$ value in each of the domains composing the $\varepsilon$ axis (see Subsec. IV.3).
@@
As mentioned above, charge accretion imbalance is guaranteed if $b_{0}^{+}<b_{1}^{-}$. From the behavior of these two quantities as presented in Fig. 6, we see that this inequality is indeed satisfied in the range $\varepsilon>\varepsilon_{\mathrm{cross}}\approx 0.289$ – implying that charge accretion is imbalanced in this entire range.
@@
Figure 6: The parameters $b_{0}^{\pm},b_{1}^{\pm}$ (which provide corresponding bounds on the absorption cross sections $\sigma^{\pm}$) as a function of the electromagnetic field strength parameter $\mathopen{}\mathclose{{\left|\varepsilon}}\right|$ along with the analytical curve representing $b_{h}$, and the universal $\varepsilon\to$-$\infty$ limit $b_{1}^{\mathrm{univ}}$, for $E=\sqrt{2},\,\alpha=0.75$. The curves for $b_{0}^{+}$ and $b_{1}^{-}$ intersect at $\varepsilon_{\mathrm{cross}}=0.289063$, above which $b_{0}^{+}<b_{1}^{-}$, demonstrating accretion imbalance. At $\varepsilon\to\infty$, $b_{0}^{+}$ and $b_{1}^{+}$ both vanish along with $b_{h}$ while the latter serves as an analytical upper bound on them. The curves for $b_{0}^{-}$ and $b_{1}^{-}$ converge towards a different constant value, which, for $b_{1}^{-}$, can be identified with $b_{1}^{\mathrm{univ}}$.
@@ VI Discussion
In this paper we have studied charge accretion into a rotating BH embedded in an asymptotically uniform magnetic field, described by Wald’s solution (Wald1974) for $A_{\mu}\mathopen{}\mathclose{{\left(x}}\right)$ (presented in Eq. (1)). We set the BH charge to the value proposed by Wald as the saturation charge for charged particle accretion, $Q_{\mathrm{w}}=2B_{0}J$. Having in mind a dilute environment where single-particle dynamics dominates, we compute upper and lower bounds on charged particles’ absorption cross section and show that at least for large enough magnetic field, the cross sections for the two types of oppositely charged particles are not equal, implying an imbalance in charge accretion at $Q=Q_{\mathrm{w}}$. This is our main result, which implies that $Q_{\mathrm{w}}$ should not be thought of as the universal model-independent saturation charge: rather, the saturation charge depends on the specifics of the accretion model. This result is somewhat reminiscent of the observation made in (Adari2023), although made in a different context and system (and a different setup than the one considered by Wald), that “dynamics upstage screening”. Nevertheless, at least in an important (and likely astrophysically relevant) region of the parameter space of the model, Wald’s charge $Q_{\mathrm{w}}$ still seems to provide a good approximation to the actual saturation charge, as we further discuss below.
For both positive and negative particle charge, there is a “central absorption disk” with cross sectional radius $b_{1}$ such that all orbits with $b<b_{1}$ fall into the BH. We solved the EOM numerically for both attracted and repelled charges and found $b_{1}$ for various choices of $\varepsilon$ (as well as $E$ and $\alpha$) providing the lower bound $\pi\mathopen{}\mathclose{{\left(b_{1}}}\right){}^{2}$ on the absorption cross section. We found that in the case of an attracted charge, $b_{1}^{-}$ tends toward a constant value of order a few times $M$ at large dimensionless field strength $\mathopen{}\mathclose{{\left|\varepsilon}}\right|$, depending only on the spin parameter $\alpha$. This was found by using the $\varepsilon\to$-$\infty$ universal limit of the EOM introduced in Subsec. III.3, where the trajectories become formally light-like and independent of $E$ (and of $\varepsilon$ obviously). This limit (which was also numerically verified; see Fig. 2) provides enhanced – yet incomplete – analytical control which improves our global understanding of the critical trajectory $b_{1}^{-}$. In contrast, for the repelled charge we find that for $\varepsilon\gg 1$ the corresponding critical trajectory converges towards the axis of symmetry $\rho=0$, and in particular, the impact parameter $b_{1}^{+}$ decreases to zero.
@@ Acknowledgements.
A. Ori would like to thank Bob Wald for helpful discussions. This work was supported in part by the Israel Science Foundation (grant No. 2047/23).
Appendix A Phases of the Critical Trajectory
While the critical trajectory is oscillatory for all choices of $E$ and $a$, we observe that there is a critical value of $\varepsilon$ which separates the space of critical trajectories into two qualitatively different phases, depending only on the value of $\varepsilon$. These two phases are characterized by an oscillation that is symmetric with respect to the equatorial plane for $\varepsilon<\varepsilon_{\mathrm{crit}}$, and an oscillation that occurs only above the equatorial plane in the case of $\varepsilon_{\mathrm{crit}}<\varepsilon$ as can be seen in Fig. 9. It is important to note that in both phases, both the $r$ and $\theta$ coordinates oscillate, as can be seen in Fig. 10. In the limit of $\varepsilon\rightarrow\varepsilon_{\mathrm{crit}}$, the critical trajectory is expected to converge towards a circular orbit on the equatorial plane.
@@ Appendix A Phases of the Critical Trajectory
Figure 9: The two different phases of the critical trajectory. The second column is a closeup of the first, focused on the region near the asymptotic locus of the precisely critical trajectory (red). The first row represents the phase $\varepsilon<\varepsilon_{\mathrm{crit}}$, where the critical trajectory oscillates symmetrically around the equatorial plane. The second row represents the phase $\varepsilon_{\mathrm{crit}}<\varepsilon$, where the critical trajectory oscillates without crossing the equatorial plane. The blue line represents an incoming (near-)critical trajectory, and the dashed lines show outgoing near-critical trajectories with $b\gtrless b_{1}$ which respectively escape to infinity (green) or fall into the BH (purple). Note that the horizontal axis representing the cylindrical radius has been extended for a better visual representation of the BH and the trajectories.
Figure 10: Oscillations of $r$ (top panel) and $\theta$ (bottom panel) as a function of the proper time $\tau$. The trajectory presented here is an infalling trajectory which is very close to the critical one, with $E=\sqrt{2},\alpha=0.5$, and $b=4.450818236305003$.
@@
In this appendix we demonstrate the fact stated in Subsec. IV.4, regarding the positivity of $n_{\mathrm{u},r}$ when evaluated at $r=r_{+},b=b_{h}$. This fact imminently leads to a forbidden region surrounding the BH, which plays a significant role in the justification to use $b_{h}$ as an upper bound on the accretion cross section of repelled particles.

### v3

@@ IV.3 Critical points and their classification
We conclude that all globally connecting and disconnecting points must correspond to solutions of this set of equations [^7] This system of equations can easily be converted to a set of polynomial equations in all its variables, by replacing the variable $\theta$ with $v\equiv\sin^{2}\theta$ – which is more convenient for analysis.. This fact provides a convenient method for finding the globally (dis)connecting points by numerically solving this set of equations. Mathematically, this set of equations typically has numerous solutions but we restrict ourselves to “physical” solutions, which are real solutions with $r>r_{+},0<\sin^{2}{\theta}\leq 1$ and $b>0$. We refer to these “physical” solutions as critical points [^8] Not to be confused with the notion of “critical trajectory” which corresponds to $b=b_{1}$ mentioned in Subsec. III.2., [^9] Interestingly, critical points can be shown to fully correspond to circular trajectories at a (not necessarily equatorial) constant $\theta$. We hope to elaborate on this elsewhere.. We empirically find that there always are at least one and at most four such critical points (for given $0<\alpha<1,\>E>1$ and $\varepsilon\neq 0$) [^10] Practically, we find the critical points by solving the polynomial variant of this set of equations (see footnote 7). To this end, we use Wolfram Mathematica’s command NSolve, which solves this algebraic system easily and with any desired precision. .
@@
Among the relevant points, some satisfy $n_{\mathrm{u},b}>0$, which implies that two allowed regions locally disconnect as $b$ is increased at this point in space; see Fig. 3. We refer to such relevant points as locally disconnecting points. Likewise, relevant points at which $n_{\mathrm{u},b}<0$, will be referred to as locally connecting points. Any globally (dis)connecting point must in particular be a locally (dis)connecting point. Importantly, note that if there is only one locally disconnecting point, it must correspond to $b_{0}$.
As mentioned above, in principle there might be “energetic rings”. In this case, there must be at least one globally connecting point – and also at least two globally disconnecting ones. As a matter of fact, based on the local analysis of critical points developed above, we conclude (by a combination of analytical considerations and empirical findings, which we hope to present elsewhere) that there are no locally connecting points for any point in the $E$ and $\alpha$ parameter space. This of course implies that there are no globally connecting points, and thus, there are no “energetic rings”. This means that the entire domain $b<b_{0}$ is energetically allowed for accretion, while the entire domain $b>b_{0}$ is forbidden.
@@
Using the latter method we find that in the specific case presented in Subsec. V.2 ($\alpha=0.75,E=\sqrt{2}$), in the first four domains there is only a single locally disconnecting point – which must be $b_{0}$. In particular, this is the situation for the specific value $\varepsilon=2$ discussed in further details in Subsec. V.1 (this particular $\varepsilon$ value belongs to the first domain). In the fifth domain, corresponding to $\varepsilon$ values larger than the last transition point $\varepsilon_{\mathrm{last}}\approx 6.8702$, there actually are two locally disconnecting points. The identification of $b_{0}$ in this case is done using a complementary analytical method as we describe in the next subsection. The implementation of the classification process for this last $\varepsilon$ domain is demonstrated in App. D and compactly summarized in Table 1.
@@ IV.4 An analytical upper bound
Considering first the situation at $b=b_{h}$, in App. B we show that at $r=r_{+}$, $n_{\mathrm{u},r}>0$ is satisfied for all $\theta$ (including $\theta=0$), which implies that $n_{\mathrm{u}}>0$ throughout some $r>r_{+}$ neighborhood of the event horizon. In other words, particles coming towards the BH with $b=b_{h}$ cannot be absorbed as their presence near the event horizon is forbidden. Because $n_{\mathrm{u}}$ is continuous (particularly in $b$), there must exist a neighborhood of $b=b_{h}$ throughout which the horizon is still surrounded by a forbidden region.
@@ V Charge Accretion (Im)balance
As explained in sections III and IV, the lower bound on the absorption cross section can be found by numerically evaluating $b_{1}$, and the upper bound by finding $b_{0}$ – or alternatively, for repelled particles, using $b_{h}$. We implemented these methods using Wolfram Mathematica to compute these two bounds. Throughout this section we focus of the specific case of $\alpha=0.75$ and $E=\sqrt{2}$. In the first subsection we specifically consider the value $\mathopen{}\mathclose{{\left|\varepsilon}}\right|=2$ and demonstrate that charge accretion is not balanced at $Q=Q_{\mathrm{w}}$. In the second subsection we extend the analysis to the entire range of $\varepsilon$ and find that charge accretion imbalance persists for all $\mathopen{}\mathclose{{\left|\varepsilon}}\right|$ values greater than a certain, relatively small, parameter $|\varepsilon|_{\mathrm{cross}}$ to be specified below.
@@ V.1 A specific example of accretion imbalance
In this subsection we focus on the specific case $|\varepsilon|=2$ (with $E=\sqrt{2}$ and $\alpha=0.75$). In this case, following the method described in Subsec. III.2, we first find that for the attracted charge $b_{1}^{-}\approx 4.216$. Then, implementing the methods described in Sec. IV for the repelled charge, we find that at $\varepsilon=2$ there is only a single critical point $(r_{\mathrm{crit}}\approx 1.706,\theta_{\mathrm{crit}}=\frac{\pi}{2},b_{\mathrm{crit}}\approx 2.427)$. A map of the allowed and forbidden regions at $b=b_{\mathrm{crit}}$ is presented in Fig. 5. As explained in Subsec. IV.2 (combined with Subsec. IV.3), in such a case, $b_{\mathrm{crit}}$ must correspond to $b_{0}$ – namely, $b_{0}^{+}\approx 2.427$. This is a relevant critical point (namely, it satisfies $\det H<0$ and $n_{\mathrm{u},rr}<0$) which is also locally disconnecting ($n_{\mathrm{u},b}>0$) – as it should.
@@
Figure 5: Map of the allowed and forbidden regions at $b=b_{0}$ for $\alpha=0.75,E=\sqrt{2}$ and $\varepsilon=2$. For this choice of $\varepsilon$ value there is only one critical point (circled in red), which necessarily corresponds to $b_{0}$. In this case, $b_{0}\approx 2.4273$.
@@
In this subsection we generalize the analysis and results presented in the previous subsection to the entire $\varepsilon$ axis. As already mentioned above (see Subsec. IV.3), to find $b_{0}$, we carry this exploration of the $\varepsilon$ axis using two independent methods (both yielding precisely the same results). The first one involves a dense sampling of $\varepsilon$ values along the $\varepsilon$ axis: we scanned the range $0<\mathopen{}\mathclose{{\left|\varepsilon}}\right|\leq 10$ with an increment $\Delta\varepsilon=0.01$, and in addition, we scanned the larger range $0<\mathopen{}\mathclose{{\left|\varepsilon}}\right|\leq 100$ with $\Delta\varepsilon=0.1$. The second method is more analytical in nature and requires the investigation of only a single representative $\varepsilon$ value in each of the domains composing the $\varepsilon$ axis (see Subsec. IV.3).
@@
As mentioned above, charge accretion imbalance is guaranteed if $b_{0}^{+}<b_{1}^{-}$. From the behavior of these two quantities as presented in Fig. 6, we see that this inequality is indeed satisfied in the range $|\varepsilon|>|\varepsilon|_{\mathrm{cross}}\approx 0.289$ – implying that charge accretion is imbalanced in this entire range.
@@
Figure 6: The parameters $b_{0}^{\pm},b_{1}^{\pm}$ (which provide corresponding bounds on the absorption cross sections $\sigma^{\pm}$) as a function of the electromagnetic field strength parameter $\mathopen{}\mathclose{{\left|\varepsilon}}\right|$ along with the analytical curve representing $b_{h}$, and the universal $\varepsilon\to$-$\infty$ limit $b_{1}^{\mathrm{univ}}$, for $E=\sqrt{2},\,\alpha=0.75$. The curves for $b_{0}^{+}$ and $b_{1}^{-}$ intersect at $|\varepsilon|_{\mathrm{cross}}\approx 0.289063$, above which $b_{0}^{+}<b_{1}^{-}$, demonstrating accretion imbalance. At $\varepsilon\to\infty$, $b_{0}^{+}$ and $b_{1}^{+}$ both vanish along with $b_{h}$ while the latter serves as an analytical upper bound on them. The curves for $b_{0}^{-}$ and $b_{1}^{-}$ converge towards different limiting values, which, for $b_{1}^{-}$, can be identified with $b_{1}^{\mathrm{univ}}$.
@@ VI Discussion
In this paper we have studied charge accretion into a rotating BH embedded in an asymptotically uniform magnetic field, described by Wald’s solution (Wald1974) for $A_{\mu}\mathopen{}\mathclose{{\left(x}}\right)$ (presented in Eq. (1)). We set the BH charge to the value proposed by Wald as the saturation charge for charged particle accretion, $Q_{\mathrm{w}}=2B_{0}J$. Having in mind a dilute environment where single-particle dynamics dominates, we compute upper and lower bounds on charged particles’ absorption cross sections and show that at least for large enough magnetic field, the cross sections for the two types of oppositely charged particles are not equal, implying an imbalance in charge accretion at $Q=Q_{\mathrm{w}}$. This is our main result, which implies that $Q_{\mathrm{w}}$ should not be thought of as the universal model-independent saturation charge: rather, the saturation charge depends on the specifics of the accretion model. This result is somewhat reminiscent of the observation made in (Adari2023), although made in a different context and system (and a different setup than the one considered by Wald), that “dynamics upstage screening”. Nevertheless, at least in an important (and likely astrophysically relevant) region of the parameter space of the model, Wald’s charge $Q_{\mathrm{w}}$ still seems to provide a good approximation to the actual saturation charge, as we further discuss below.
For both positive and negative particle charge, there is a “central absorption disk” with cross sectional radius $b_{1}$ such that all orbits with $b<b_{1}$ fall into the BH. We solved the EOM numerically for both attracted and repelled charges and found $b_{1}$ for various choices of $\varepsilon$ (as well as $E$ and $\alpha$) providing the lower bound $\pi\mathopen{}\mathclose{{\left(b_{1}}}\right){}^{2}$ on the absorption cross section. We found that in the case of an attracted charge, $b_{1}^{-}$ tends toward a constant value of order a few times $M$ at large dimensionless field strength $\mathopen{}\mathclose{{\left|\varepsilon}}\right|$, depending only on the spin parameter $\alpha$. This was found by using the $\varepsilon\to$-$\infty$ universal limit of the EOM introduced in Subsec. III.3, where the trajectories become formally light-like and independent of $E$ (and, obviously, of $\varepsilon$). This limit (which was also numerically verified; see Fig. 2) provides enhanced – yet incomplete – analytical control which improves our global understanding of the critical trajectory $b_{1}^{-}$. In contrast, for the repelled charge we find that for $\varepsilon\gg 1$ the corresponding critical trajectory converges towards the axis of symmetry $\rho=0$, and in particular, the impact parameter $b_{1}^{+}$ decreases to zero.
@@ Acknowledgements.
We are grateful to Noa Zilberman for helpful comments. A. Ori would like to thank Bob Wald for helpful discussions. This work was supported in part by the Israel Science Foundation (grant No. 2047/23).
Appendix A Phases of the Critical Trajectory $b_{1}$
While the critical trajectory $b=b_{1}$ is oscillatory for all choices of $E$ and $a$, we observe that there is a critical value of $\varepsilon$ which separates the space of critical trajectories into two qualitatively different phases, depending only on the value of $\varepsilon$. These two phases are characterized by an oscillation that is symmetric with respect to the equatorial plane for $\varepsilon<\varepsilon_{\mathrm{crit}}$, and an oscillation that occurs only above the equatorial plane in the case $\varepsilon>\varepsilon_{\mathrm{crit}}$, as can be seen in Fig. 9. It is important to note that in both phases, both the $r$ and $\theta$ coordinates oscillate, as can be seen in Fig. 10. In the limit of $\varepsilon\rightarrow\varepsilon_{\mathrm{crit}}$, the critical trajectory is expected to converge towards a circular orbit on the equatorial plane.
@@ Appendix A Phases of the Critical Trajectory $b_{1}$
Figure 9: The two different phases of the critical trajectory. The second column is a closeup of the first, focused on the neighborhood of the asymptotic oscillatory trajectory, further explained below. The first row represents the phase $\varepsilon<\varepsilon_{\mathrm{crit}}$, where the critical trajectory oscillates, asymptotically symmetrically, around the equatorial plane. The second row represents the phase $\varepsilon>\varepsilon_{\mathrm{crit}}$, where the critical trajectory oscillates without crossing the equatorial plane. The solid blue line represents the incoming critical trajectory $b=b_{1}$, which at late times asymptotically approaches a precisely periodic trajectory (the solid red curve). The two dashed lines show near-critical orbits coming from infinity with $b\gtrless b_{1}$ along trajectories very close to the blue one, which either escape to infinity (green) or fall into the BH (purple) after oscillating several times near the red curve.
Figure 10: Oscillations of $r$ (top panel) and $\theta$ (bottom panel) as a function of the proper time $\tau$. The trajectory presented here is an infalling trajectory which is very close to the critical one, with $E=\sqrt{2},\alpha=0.5,\varepsilon=-1$, and $b=4.450818236305003$.
@@
In this appendix we demonstrate the fact stated in Subsec. IV.4, regarding the positivity of $n_{\mathrm{u},r}$ when evaluated at $r=r_{+},b=b_{h}$. This implies the existence of a forbidden region surrounding the BH, which is central to justifying the use of $\pi(b_{h})^{2}$ as an upper bound on the accretion cross section for repelled particles.